A rectangle has a length twice it's width, and area 288cm squared, what is the perimeter?
I have tried square rooting but I don't know if it's right.

Answer by Cromlix(4381)

Hi there,
Make the width = 'x'
Length then = 2x
Area = Width x Length
288 = x(2x)
2x^2 = 288
x^2 = 144
x = 12
Width = 12 cm (x)
Length = 24 cm (2x)
Perimeter = 2(Width) + 2(Length)
Perimeter = 2(12) + 2(24)
Perimeter = 24 + 48
Perimeter = 72 cm.
Hope this helps :-)